### v1

# Minimal commutant and double commutant property for analytic Toeplitz operators

Abstract
In this paper we study the minimality of the commutant of an analytic Toeplitz operator $M_{\varphi}$, when $M_{\varphi}$ is defined on the Hardy space $H^{2}(\mathbb{D})$ and $\varphi\in H^{\infty}(\mathbb{D})$, denotes a bounded analytic function on $\mathbb{D}$. Specifically we show that the commutant of $M_{\varphi}$ is minimal if and only if the polynomials on $\varphi$ are weak-star dense in $H^{\infty}(\mathbb{D})$, that is, $\varphi$ is a weak-star generator of $H^{\infty}(\mathbb{D})$. We use our result to characterize when the double commutant of an analytic Toeplitz operator $M_{\varphi}$ is minimal, for a large class of symbols $\varphi$. Namelly, when $\varphi$ is an entire function, or more generally when $\varphi$ belongs to the Thomson-Cowen’s class.
keywords:
Analytic Toeplitz operator, Minimal commutant , Double commutant property.
MSC:
47B35, 47B38, 30J05
1 Introduction
Let $\mathcal{L}(H)$ stand for the algebra of all bounded linear operators on a Hilbert space $H$. If $T\in\mathcal{L}(H)$ then the following algebras containing $T$ arise naturally on $\mathcal{L}(H)$: $\overline{\textrm{Alg}(T)}^{WOT}$, $\{T\}^{\prime}$ and $\{T\}^{\prime\prime}$. When $\textrm{Alg}(T)$ denotes the unital algebra generated by $T$, and $\overline{\textrm{Alg}(T)}^{WOT}$ denotes its closure in the weak operator topology. The commutant of $T$, denoted by $\{T\}^{\prime}$, is the set of all operator $A$, such that $AT=TA$. And the double commutant of $T$, denoted by $\{T\}^{\prime\prime}$, is the set of all operators $A$ such that $AB=BA$ for all $B\in\{T\}^{\prime}$.
In general, the following inclusions $\textrm{Alg}(T)\subset\{T\}^{\prime\prime}\subset\{T\}^{\prime}$ are trivial, and since the algebras $\{T\}^{\prime},\{T\}^{\prime\prime}$ are closed in the weak operator topology, we have also $\overline{\textrm{Alg}(T)}^{WOT}\subset\{T\}^{\prime\prime}\subset\{T\}^{\prime}$. The commutant and double commutant of an operator determine it structure. Therefore it is important to know as much as possible about the commutant of given operator. For instance, it is natural to wonder when $\overline{\textrm{Alg}(T)}^{WOT}=\{T\}^{\prime}$, that is, when $T$ has the minimal commutant property, because $T$ enjoys this property if and only if the invariant and hyperinvariant subspace lattices of $T$ coincide. Also, $T$ is said to have the double commutant property if its double commutant is equal to the weak closure of the polynomials on $T$.
The literature on these topics is very big. On general operator theory, we cannot talk about double commutant without citing J. Von Neumann’s double commutant theorem [30]. Herrero [17] was trying to characterize the norm closure of the class of operators having the minimal commutant property. This question, that still resists, firstly appeared at the end of the Ph’D thesis by Turner ([29]), but also as Herrero point out, it was formulated in private communication by J. B. Conway.
On Concrete Operator Theory, to solve questions concerning commutants of analytic Toeplitz operators was a mainstream of research in the early 1970s. The main ancestors of this paper are results of Abrahamse, Baker, Cowen, Deddens, Shields, Thomson and many others ([9, 14, 24, 27, 29]). The study of commutants and double commutants of composition operators is more recent. The first findings appeared in the 90’s [7, 8, 32]. More recently, composition operators, induced by linear fractional selfmaps of the unit disk, having the minimal commutant property [18] and the double commutant property [19] have been characterized.
In the PhD thesis of Turner ([29]) was formulated the following question
Question 1: For which maps $\varphi\in H^{\infty}(\mathbb{D})$ the analytic Toeplitz operator $M_{\varphi}$ has the double commutant property?
As we will see in the present paper, Question 1, is related to the minimal commutant property for analytic Toeplitz operator. Thus the following question arises:
Question 2: For which maps $\varphi\in H^{\infty}(\mathbb{D})$ the analytic Toeplitz operator $M_{\varphi}$ has the minimal commutant property?
In this paper we show that an analytic Toeplitz operator $M_{\varphi}$ has a minimal commutant if and only if its inducing map $\varphi$ is a weak-star generator of $H^{\infty}(\mathbb{D})$. Recall that a map $\varphi\in H^{\infty}(\mathbb{D})$ is a weak star generator of $H^{\infty}(\mathbb{D})$ if and only if the polynomials on $\varphi$ are weak-star dense in $H^{\infty}(\mathbb{D})$ ([22, 23]). The reason way of this phenomenon is that the closure of linear subspaces on $H^{\infty}(\mathbb{D})$ on some different topologies coincide. The proof use the Banach and Mazurkiewicz’s sequential approach of weak-star topology ([4, Annexe]).
As a consequence of this Toeplitz minimal commutant theorem we could characterize when an univalent symbol $\varphi$ induces an Toeplitz operator with the double commutant property. Moreover, with the help of the Baker-Deddens-Ulmann’s factorization theorem [3] we will characterize the entire functions $f$ that induces an analytic Toeplitz operator $M_{f}$ with the double commutant property. The result can be also extended for analytic functions $\varphi$ in the Thomson-Cowen’s class, that is, the functions $\varphi\in H^{\infty}(\mathbb{D})$ such that, for some $a\in\mathbb{D}$ the inner part of $\varphi(z)-\varphi(a)$ is a finite Blasckhe product. The Thomson-Cowen’s class includes the non-constant elements in $H^{\infty}(\overline{\mathbb{D}})$.
[…]
The paper is structured as follows. In Section 2 we show that an analytic Toeplitz operator $M_{\varphi}$ has the minimal commutant property if and only if $\varphi$ is a weak star generator of $H^{\infty}(\mathbb{D})$. As a first consequence, we will characterize the univalent symbols $\varphi$ such that $M_{\varphi}$ has the double commutant property. We present these results by recalling some classical ideas and motivating the reader with concrete examples when it is possible.
In Section 3 we focus our attention on symbols $\varphi\in H^{\infty}(\mathbb{D})$ that single cover an open neighbourhood and on a classical result by Deddens and Wong (we include a proof that uses only function theory) that determines the commutant of $M_{\varphi}$. These maps $\varphi$ will give us the first examples of operators $M_{\varphi}$ that don’t have the double commutant property.
In Section 4 we discover the first examples of non-univalent functions $\varphi\in H^{\infty}(\mathbb{D})$ that induce $M_{\varphi}$ with the double commutant property. For instance when $\varphi$ is an inner function. Although these results follow from a more general result by Turner that ensures that every non-unitary isometry has the double commutant property, we will show an alternative proof. Finally, in Section 5 we characterize when an entire function or a function in the Thomson-Cowen class has the double commutant property. The paper closes with a brief section with concluding remarks and open questions.
2 Analytic Toeplitz operators with a minimal commutant
The pioneering results on commutants of Toeplitz operators date back to the work of Shields and Wallen [24]. They consider commutants of operators that can be seeing as multiplication by $z$ in a Hilbert space of analytic functions. In fact, it was noticed later by Deddens and Wong [14] that their method can be used to show that if $\varphi\in H^{\infty}(\mathbb{D})$ is univalent then $\{M_{\varphi}\}^{\prime}=\{M_{h}\,:\,h\in H^{\infty}(\mathbb{D}\}$. That is, if $\varphi\in H^{\infty}(\mathbb{D})$ is univalent then $\{M_{\varphi}\}^{\prime}=\{M_{z}\}^{\prime}$. Moreover, they show that if $T\in\{M_{\varphi}\}^{\prime}$ then $T$ can be approximate in the weak operator topology by polynomials on $M_{z}$.
However, in general we cannot assert that if $\varphi$ is univalent then each $T\in\{M_{\varphi}\}^{\prime}$ can be approximate in the weak operator topology by polynomials on $M_{\varphi}$, that is, we cannot assert that if $\varphi$ is univalent then $M_{\varphi}$ has a minimal commutant. To assert minimality on the commutant of $M_{\varphi}$ we need an extra hyphotesis. Specifically the polynomials on $\varphi$ must be weak-star dense in the space $H^{\infty}(\mathbb{D})$. In fact, we will show that $M_{\varphi}$ has a minimal commutant if and only if the polynomials on $\varphi$ are weak-star dense in $H^{\infty}(\mathbb{D})$.
As a byproduct of the Toeplitz minimal commutant’s theorem we characterize when a Toeplitz operator $M_{\varphi}$ with univalent symbol $\varphi$ has the double commutant property.
The univalence hypothesis is very important in this circle of ideas. For instance, in this theorem due to Shields and Wallen the univalence is used in part (a). We include a proof for the sake of complementness. Let us denote by $k_{a}(z)$ the Hardy reproducing kernel.
Theorem 2.1 (Shields-Wallen).
Assume that $\varphi\in H^{\infty}(\mathbb{D})$ is univalent.
If $T\in\{M_{\varphi}\}^{\prime}$ then there exists $h\in H^{\infty}(\mathbb{D})$ such that $T=M_{h}$.
For each $T\in\{M_{\varphi}\}^{\prime}$ there exists a sequence of polynomials $p_{n}$ such that $p_{n}(M_{z})\to T$ in the weak operator topology.
Proof.
Indeed, if $T\in\{M_{\varphi}\}^{\prime}$ then $T^{\star}M_{\varphi}^{\star}=M_{\varphi}^{\star}T^{\star}$. By observing that $M_{\varphi}^{\star}k_{a}=\overline{\varphi(a)}k_{a}$ we get
$M_{\varphi}^{\star}T^{\star}k_{a}=\overline{\varphi(a)}T^{\star}k_{a}.$
Since $\textrm{ker}(M_{\varphi}^{\star}-\overline{\varphi(a)}I)$ is one dimensional, we get that there exists some complex number $\overline{h(a)}$ such that $T^{\star}k_{a}=\overline{h(a)}k_{a}$ for each $a\in\mathbb{D}$.
Since the function $a\in\mathbb{D}\to k_{a}\in H^{2}(\mathbb{D})$ is analytic on $\mathbb{D}$ we get that $h$ is also analytic on $\mathbb{D}$. Moreover, since $T^{\star}$ is bounded we get that $h\in H^{\infty}(\mathbb{D})$.
Finally, we consider the analytic Toeplitz operator $M_{h}$, and we see that $T^{\star}k_{a}=M_{h}^{\star}k_{a}$ for all $a\in\mathbb{D}$. Since $\{k_{a}(z)\,:\,a\in\mathbb{D}\}$ is a total set in $H^{2}(\mathbb{D})$ we obtain that $T=M_{h}$ which proves (a).
Set $T\in\{M_{\varphi}\}^{\prime}$, since $\varphi$ is univalent there exists $h\in H^{\infty}(\mathbb{D})$ such that $T=M_{h}$. For $h\in H^{\infty}(\mathbb{D})$ we consider $(p_{n}(z))$ the Féjèr polynomials of $h$. Hence $p_{n}(z)\to h(z)$ uniformly on compact subsets of $\mathbb{D}$ and $\|p_{n}\|_{\infty}\leq\|h\|_{\infty}$, $n\geq 1$. Therefore, for each $a\in\mathbb{D}$ we obtain:
$\langle p_{n}(M_{z})f,k_{a}\rangle=p_{n}(a)f(a)\to h(a)f(a)=\langle M_{h}f,k_{a}\rangle.$
The above equality is also true for linear finite combinations of the set $D=\{k_{a}\,:\,a\in\mathbb{D}\}$. Since $D$ is a total set, given $g\in H^{2}(\mathbb{D})$ there exist a sequence $g_{l}$ of linear combinations of the set $D$ such that $\|g-g_{l}\|_{2}\to 0$.
We have that $\|p_{n}(M_{z})\|=\|p_{n}\|_{\infty}\leq\|h\|_{\infty}$, therefore:
$\lim_{n}\langle p_{n}(M_{z})f,g\rangle=\lim_{n}\lim_{l}\langle p_{n}(M_{z})f,g_{l}\rangle=\lim_{l}\langle M_{h}f,g_{l}\rangle=\langle M_{h}f,g\rangle,$
that is, $p_{n}(M_{z})\to M_{h}$ in the weak operator topology, as we desired. ∎
We wish to characterize the analytic Toeplitz operators $M_{\varphi}$ with a minimal commutant. The next results show that univalence on $\varphi$ is a necessary condition.
Proposition 2.2.
If $M_{\varphi}$ has a minimal commutant then $\varphi$ must be univalent.
Proof.
Indeed, if $\varphi$ is not univalent, there exists two different points in $\mathbb{D}$, $a$ and $b$, such that $\varphi(a)=\varphi(b)=c$.
Recall that an operator $T$ has minimal commutant if and only if $T^{\star}$ has minimal commutant. We will conclude if we prove that $M_{\varphi}^{\star}$ has not a minimal commutant. To show that, we will see that $M_{z}^{\star}\notin\overline{\textrm{Alg}(M_{\varphi}^{\star})}^{WOT}$.
Firstly, let us observe that $\overline{\textrm{Alg}(M_{\varphi}^{\star})}^{WOT}=\overline{\textrm{Alg}(M_{\varphi}^{\star}-cI)}^{WOT}=\overline{\textrm{Alg}(M_{\varphi-c}^{\star})}^{WOT}$. In particular for $c=\varphi(a)$. Set $\phi=\varphi-\varphi(a)$.
By way of contradiction, let us assume that there exists a net $p_{d}(M_{\phi}^{\star})$, $d\in D$ such that $p_{d}(M_{\phi}^{\star})\rightarrow M_{z}^{\star}$. Let us denote by $K_{a}$ is the reproducing kernel at $a$, normalized. Since $M_{\phi}^{\star}(K_{a})=M_{\phi}^{\star}(K_{b})=0$ then
$\lim_{d\in D}\langle p_{d}(M_{\phi}^{\star})K_{a},K_{a}\rangle=\lim_{d\in D}p_{d}(0)=\langle M_{z}^{\star}K_{a},K_{a}\rangle=\overline{a}\cdot\|K_{a}\|^{2}=\overline{a}.$
On the other hand
$\lim_{d\in D}\langle p_{d}(M_{\phi}^{\star})K_{b},K_{b}\rangle=\lim_{d\in D}p_{d}(0)=\langle M_{z}^{\star}K_{b},K_{b}\rangle=\overline{b},$
that is $\lim_{d}p_{d}(0)=\overline{a}\neq\overline{b}=\lim_{d}p_{d}(0)$, a contradiction. ∎
The space $H^{\infty}(\mathbb{D})$ is a closed subspace of $L^{\infty}(\partial\mathbb{D})$. Thus $H^{\infty}(\mathbb{D})$ can be seen as the dual of a quotient of $L^{1}(\partial\mathbb{D})$. Since $H^{\infty}(\mathbb{D})$ is the dual of a separable Banach space, it is possible to define a weak-star topology in $H^{\infty}(\mathbb{D})$ (see [23]). Let us denote this topology by $\tau_{\star}$.
A function $\varphi\in H^{\infty}(\mathbb{D})$ is said to be a weak-star generator of $H^{\infty}(\mathbb{D})$ provided the polynomials on $\varphi$ are weak-star dense in $H^{\infty}(\mathbb{D})$.
The main result of this section requires some technicalities on some topologies defined on $H^{\infty}(\mathbb{D})$. Asssume that $\varphi$ is univalent. By applying Proposition 2.2, if $M_{h_{\alpha}}$ is a net converging to an operator $T$ in the weak operator topology then $T=M_{h}$ for some $h\in H^{\infty}(\mathbb{D})$. On $H^{\infty}(\mathbb{D})$ we can consider also the relative topology induced by weak operator topology. If $g\in H^{\infty}(\mathbb{D})$ a basic neighbourhood of $g$ is defined as:
$V_{wot}(g;f_{1},f_{2};\varepsilon)=\{h\in H^{\infty}(\mathbb{D})\,:\,|\langle(M_{h}-M_{g})f_{1},f_{2}\rangle|<\varepsilon\}$
for some $f_{1},f_{2}\in H^{2}(\mathbb{D})$. Let us denote this topology by $\tau_{wot}$.
Moreover, if $\varphi$ is univalent then by Proposition 2.2 the WOT closure of the polynomials on $M_{\varphi}$ is a subspace of $\{M_{z}\}^{\prime}$. Thus it can be identified as a subspace of $H^{\infty}(\mathbb{D})$ which is exactly the $\tau_{wot}$-closure of the polynomials on $\varphi$.
Analogously, we can make the same observation with a stronger topology on $\mathcal{L}(H^{2}(\mathbb{D}))$: the $\sigma$-weak star or ultraweak star operator topology. Let us denote this topology by $\tau_{\sigma*}$. If $g\in H^{\infty}(\mathbb{D})$ then a basic neighbourhood of $g$ on $\tau_{\sigma*}$ is defined as:
$V_{\sigma*}(g;(f_{n})_{n},(g_{n})_{n};\varepsilon)=\left\{h\in H^{\infty}\,:\,\left|\sum_{n=1}^{\infty}\langle(M_{g}-M_{h})f_{n},g_{n}\rangle\right|<\varepsilon\right\}$
[…]
Lemma 2.3.
The weak operator topology and the $\sigma$-weak operator topology restricted to $H^{\infty}(\mathbb{D})$ coincide.
Proof.
In general the $\sigma$-weak operator topology is stronger than the weak operator topology. Hence, it is sufficient to show that each $\tau_{\sigma*}$ basic neighborhoood of 00 is is a $\tau_{wot}$ neigbourhood of 00. A basic $\tau_{\sigma*}$ neighborhood of 00 is defined as:
$V_{\sigma*}(0;(f_{n}),(g_{n});\varepsilon)=\left\{h\in H^{\infty}\,:\,\left|\sum_{n=1}^{\infty}\langle hf_{n},g_{n}\rangle\right|<\varepsilon\right\}.$
where $\sum_{n=1}^{\infty}\|f_{n}\|^{2}<\infty$ and $\sum_{n=1}^{\infty}\|g_{n}\|^{2}<\infty$.
If $\|\cdot\|_{1}$ denotes the standard norm in $H^{1}(\mathbb{D})$ then by applying Holder inequalities we get:
$\left\|\sum_{n=1}^{\infty}f_{n}\overline{g_{n}}\right\|_{1}$$\leq$ $\sum_{n=1}^{\infty}\|f_{n}\overline{g_{n}}\|_{1}$
$\leq$ $\sum_{n=1}^{\infty}\|f_{n}\|^{2}\|g_{n}\|^{2}$
$\leq$ $\left(\sum_{n=1}^{\infty}\|f_{n}\|^{2}\right)^{1/2}\left(\sum_{n=1}^{\infty}\|g_{n}\|^{2}\right)^{1/2}<\infty.$
Which implies that $F(z)=\sum_{n=1}^{\infty}f_{n}\overline{g_{n}}\in H^{1}(\mathbb{D})$ and by Fubini’s theorem we can change the sum and the integral:
[…]
Le us denote by $B$ a Blaschke product with the same zeros of $F$. Since $g(z)=F(z)/B(z)$ has no zeros, we can write $F(z)=B(z)g(z)^{1/2}g(z)^{1/2}$. And since $F\in H^{1}(\mathbb{D})$ we have that $B(z)g(z)^{1/2}$ and $g(z)^{1/2}$ are functions on $H^{2}(\mathbb{D})$. Therefore:
[…]
Hence,
$V_{\sigma*}(0,(f_{n}),(g_{n});\varepsilon)=V_{wot}(0;Bg^{1/2},g^{1/2};\varepsilon)$
which yields the desired result. ∎
Lemma 2.4.
Assume that $h\in H^{\infty}(\mathbb{D})$. The following conditions are equivalents:
a)
$h\in H^{\infty}(\mathbb{D})$ is weak-star limit of $p_{n}(\varphi)$.
b)
$h\in H^{\infty}(\mathbb{D})$ is if the WOT-limit of $p_{n}(M_{\varphi})$.
c)
$h\in H^{\infty}(\mathbb{D})$ is $\sigma$-weak star limit of $p_{n}(M_{\varphi})$.
Proof.
The equivalence between b) and c) follows by using Lemma 2.3. To stablish a) implies b), assume that $h$ is weak-star limit of polynomials on $\varphi$ then ([23, Lemma 1]) there exists a sequence of polynomials $p_{n}$, such that, $p_{n}(\varphi)(a)\to h(a)$ for any $a\in\mathbb{D}$ and $\|p_{n}(\varphi)\|_{\infty}<M$ for all $n$. Given $f,g\in H^{2}(\mathbb{D})$ and $\varepsilon>0$ we wish to to prove that there exists $n_{0}$ such that $|\langle(M_{p_{n}(\varphi)}-M_{h})f,g\rangle|<\varepsilon$ for all $n\geq n_{0}$.
Since the set of reproducing kernels $\{k_{a}(z)\,:\,a\in\mathbb{D}\}$ is a total set in $H^{2}(\mathbb{D})$, there exists a finite linear combinations of reproducing kernels $g_{\varepsilon}$ such that
$\|g-g_{\varepsilon}\|\leq\frac{\varepsilon}{2\|f\|(M+\|h\|_{\infty})}.$ (1)
The pointwise convergence of the hypothesis implies that $\langle(M_{p_{n}(\varphi)}-M_{h})f,k_{a}\rangle$ converges to zero. Therefore, there exists $n_{0}$ such that for $n\geq n_{0}$
$|\langle(M_{p_{n}(\varphi)}-M_{h})f,g_{\varepsilon}\rangle|<\varepsilon/2.$ (2)
Finally, by using (1) and (2) and the triangular inequality, we get that for any $n\geq n_{0}$
$|\langle(M_{p_{n}(\varphi)}-M_{h})f,g\rangle|$$\leq$ $|\langle(M_{p_{n}(\varphi)}-M_{h})f,g_{\varepsilon}\rangle|+|\langle(M_{p_{n}(\varphi)}-M_{h})f,g-g_{\varepsilon}\rangle|$
$\leq$ $\varepsilon/2+\frac{\varepsilon}{2}\frac{\|p_{n}(\varphi)-h\|_{\infty}}{M+\|h\|_{\infty}}\leq\varepsilon.$
Conversely if $M_{p_{n}(\varphi)}\to M_{h}$ in WOT then $\langle p_{n}(\varphi),k_{a}\rangle=\varphi_{n}(a)\to h(a)=\langle h,k_{a}\rangle$. On the other hand, if $\langle M_{p_{n}(\varphi)}f,g\rangle\to\langle M_{h}f,g\rangle$ for any $f,g\in H^{2}(\mathbb{D})$ then $M_{p_{n}(\varphi)}f\to M_{h}f$ weakly. Thus $M_{p_{n}(\varphi)}f$ is pointwise bounded for any $f\in H^{2}(\mathbb{D})$. Therefore by the Uniform Boundedness Principle we get that $\|M_{p_{n}(\varphi)}\|=\|p_{n}(\varphi)\|_{\infty}<M$ for all $n$. That is, $p_{n}(\varphi)\to h$ in the weak-star topology, which establish the equivalence between a) and b). ∎
Remark 2.5.
The proof of the above result gives us a little more. Specifically, on $H^{\infty}(\mathbb{D})$ the three topologies have the same convergent sequences. That is, $\psi_{n}\in H^{\infty}(\mathbb{D})$ converges weak-star to $h$ if and only if $M_{\psi_{n}}$ converges in the weak operator topology to $M_{h}$.
To show Theorem 2.6 we will use the sequential approach of the weak-star topology introduced by Banach and Mazurkiewicz ([4, Annexe]). If $\mathcal{P}$ is a linear subspace of a Banach space $X$ with separable predual, then the weak-star closure of $\mathcal{P}$ can be recovered by taking limits on $\mathcal{P}$ in a transfinite way. Specifically, if $B$ is the unit ball of $X$, the derived set of $\mathcal{P}$ is defined as:
$\mathcal{P}^{(1)}=\bigcup_{n=1}^{\infty}\overline{\mathcal{P}\cap nB}^{*},$
which represent the weak star limit of sequences on $\mathcal{P}$. In general it is well know that $\mathcal{P}^{(1)}$ do no coincide with the weak star closure of $\mathcal{P}$. In this way, it is natural to introduce derived sets for any ordinal number as follows: if $\mathcal{P}^{(\alpha)}$ has been already defined, then $\mathcal{P}^{(\alpha+1)}=(\mathcal{P}^{(\alpha)})^{(1)}$.
When the predual of $X$ is separable then there exists an ordinal $\beta$ such that $\mathcal{P}^{(\beta)}=\mathcal{P}^{(\beta+1)}=\overline{\mathcal{P}}^{*}$. That is, we can recover the weak-star closure of a subspace taking derived sets in a transfinite way. Now we are in position to show the main result of this section.
Theorem 2.6.
Set $\varphi\in H^{\infty}(\mathbb{D})$. $M_{\varphi}$ has a minimal commutant if and only if $\varphi$ is a weak-star generator of $H^{\infty}(\mathbb{D})$.
Proof.
Firstly, let us observe that if $M_{\varphi}$ has a minimal commutant then $\varphi$ is univalent (see Proposition 2.2). On the other hand if $\varphi$ is a weak star generator of $H^{\infty}(\mathbb{D})$ then also $\varphi$ must be univalent (see [22, Proposition 3]). Thus, we can suppose without loss that $\varphi$ is univalent.
Let us denote by $\mathcal{P}$ the linear subspace of the polynomials on $\varphi$. We claim that the weak-star closure of $\mathcal{P}$: $\overline{\mathcal{P}}^{*}$, and the WOT-closure of $\mathcal{P}$: $\overline{\mathcal{P}}^{WOT}$ is the same set.
Indeed, let us denote by $\mathcal{P}_{*}^{(\alpha)},\mathcal{P}_{\sigma*}^{(\alpha)},$ and $\mathcal{P}_{wot}^{(\alpha)}$ the derived set of order $\alpha$ in the weak-star, $\sigma$-weak star and WOT topologies respectively. By applying Lemma 2.4 and Remark 2.5, we get that for any ordinal $\alpha$ we obtain
$\mathcal{P}_{*}^{(\alpha)}=\mathcal{P}_{\sigma*}^{(\alpha)}=\mathcal{P}_{wot}^{(\alpha)}.$
Since the predual of $H^{\infty}$ (a quotient of $L^{1}(\mathbb{D})$) and the predual of $\mathcal{L}(H^{2}(\mathbb{D}))$ (the trace class) are separable, there exists an ordinal $\beta$ such that
$\overline{\mathcal{P}}^{\sigma*}=\mathcal{P}_{\sigma*}^{(\beta)}=\mathcal{P}_{*}^{(\beta)}=\overline{\mathcal{P}}^{*}.$
Finally, by applying Lemma 2.3 we get that $\overline{\mathcal{P}}^{\sigma*}=\overline{\mathcal{P}}^{wot}$. Now, the result follows as an observation. If $\varphi$ is univalent then by Shields-Wallen the commutant of $M_{\varphi}$ is $H^{\infty}(\mathbb{D})$. Thus, the WOT-closure of the polynomials on $\varphi$ is $H^{\infty}(\mathbb{D})$ is and only if the weak-star closure of polynomials on $\varphi$ is $H^{\infty}(\mathbb{D})$, that is, $\varphi$ is a weak-star generator of $H^{\infty}(\mathbb{D})$. ∎
Remark 2.7.
For future references, from the proof of the above result, we can deduce that if $\mathcal{M}\subset H^{\infty}(\mathbb{D})$ is a linear subspace then ¯=¯$\overline{\mathcal{M}}^{*}=\overline{\mathcal{M}}^{WOT}$.
The next result, with the help of Theorem 2.6, characterizes those univalent symbols $\varphi$ for which the analytic Toeplitz operator $M_{\varphi}$ has the double commutant property.
Corollary 2.8.
Assume that $\varphi\in H^{\infty}(\mathbb{D})$ is univalent. Then $M_{\varphi}$ has the double commutant property if and only if the polynomials on $\varphi$ are weak-star dense in $H^{\infty}(\mathbb{D})$.
Proof.
Indeed, clearly if the polynomials on $\varphi$ are weak star dense in $H^{\infty}(\mathbb{D})$ then $M_{\varphi}$ has a minimal commutant. Therefore $M_{\varphi}$ has trivially the double commutant property.
Conversely if $\varphi$ is univalent then $\{M_{\varphi}\}^{\prime}=\{M_{z}\}^{\prime}$, therefore $\{M_{\varphi}\}^{\prime\prime}=\{M_{z}\}^{\prime\prime}=\{M_{h}\,:\,h\in H^{\infty}(\mathbb{D})\}$. Therefore, since $M_{\varphi}$ has the double commutant property then
$\overline{\textrm{Alg}(M_{\varphi})}^{WOT}=\{M_{\varphi}\}^{\prime\prime}=\{M_{z}\}^{\prime}=\{M_{\varphi}\}^{\prime}.$
Thus $M_{\varphi}$ has a minimal commutant, therefore $\varphi$ is a weak-star generator of $H^{\infty}(\mathbb{D})$ as we desired. ∎
Let us remark that the minimal commutant property for an analytic Toeplitz operator $M_{\varphi}$ depends of the shape of $G=\varphi(\mathbb{D})$. For instance using several classical results on approximation we can find several examples of univalent maps $\varphi$ which are generators of $H^{\infty}(\mathbb{D})$. For instance if $\varphi(z)=z$ then the Féjèr polynomials are weak star dense on $H^{\infty}(\mathbb{D})$. If $\varphi$ maps $\mathbb{D}$ univalently onto $G$ and $G$ is a simply connected domain whose boundary is a Jordan curve, then Walsh’s Theorem asserts that the polynomials on $\varphi$ are weak star dense on $H^{\infty}(\mathbb{D})$.
As it was point out by D. Sarason [23] the domains for which $\varphi$ is a weak-star generator of $H^{\infty}$ or order 1, were characterized by Farrel in [15, 16]. Namely $\varphi$ is a weak-star generator of $H^{\infty}(\mathbb{D})$ if and only if $\varphi(\mathbb{D})$ is a Caratheodory domain. In [23] Sarason provided examples of maps $\varphi$ which are weak-star generators of $H^{\infty}(\mathbb{D})$ of order (finite) different of $1$.
On the other hand, it is simple to find maps which are not generators of $H^{\infty}(\mathbb{D})$. By [22, Proposition 1], if $\varphi$ is a generator of $H^{\infty}(\mathbb{D})$ then $\varphi$ must be univalent almost everywhere on $\partial\mathbb{D}$. Thus, if $\varphi$ maps $\mathbb{D}$ univalently onto the slit disk $\mathbb{D}\setminus[0,1)$ then $M_{\varphi}$ don’t have a minimal commutant, and also by applying Corollary 2.8 $M_{\varphi}$ don’t have the double commutant property.
Moreover, it is possible to find maps which are univalent in the unit circle for except two point which don’t induce analytic Toeplitz operators with a minimal commutant. For instance, let us consider the region formed by the disk with center zero and radius one, minus the closure of the disk with center $1/2$ and radius $1/2$. This region a like a special lunar eclipse. If we consider $\varphi$ an analytic function which maps $\mathbb{D}$ onto $G$, clearly there are bounded analytic functions on $G$ which cannot be approximate by polynomials, therefore the operator $M_{\varphi}$ has not a minimal commutant.
Figure 1: The lunar eclipse
Given $\varphi\in H^{\infty}(\mathbb{D})$, since the commutant $\{M_{\varphi}\}^{\prime}=\{M_{\varphi/R}\}^{\prime}$, we can suppose without loss of generality that $\|\varphi\|_{\infty}<1$.
From the results of Deddens ([11, 12]) there exists an important connection between analytic Toeplitz operators and composition operators.
The results obtaining here have similar flavor as the results relating to the cyclicity and the denseness of the rank of a composition operator For instance, the univalence hypothesis on the map is necessary (even the univalence almost everywhere on the unit circle $\partial\mathbb{D}$). However, there is not equivalence between cyclicity of $C_{\varphi}$ and the minimal commutant property of the operator $M_{\varphi}$. Indeed, as it was pointed out in [5] the composition operator $C_{\varphi}$ induced by $z/(2-z)$ is not cyclic, however $M_{\varphi}$ has a minimal commutant.
The rank of $C_{\varphi}$ is dense means exactly that the polynomials on $\varphi$ are dense in $H^{2}(\mathbb{D})$. If $\varphi$ is a weak-star generator of $H^{\infty}(\mathbb{D})$ then the rank of $C_{\varphi}$ is dense (see also [21]). However the converse is not true, there are maps $\varphi$ such that the polynomials on $\varphi$ are dense in $H^{2}(\mathbb{D})$ but $\varphi$ is not a weak-star generator of $H^{\infty}(\mathbb{D})$ (see [2]).
3 Double commutant property for analytic Toeplitz operators induced by maps that single cover an open neighbourhood
Maps very close to univalent ones are those $\varphi$ non-univalent for which $\varphi$ single covers an open neighbourhood $W$ of $\varphi(\mathbb{D})$. Let us denote by $S\subset H^{\infty}(\mathbb{D})$ the set of such maps. The objective of this section is to prove that if $\varphi\in S$ then $M_{\varphi}$ don’t have the double commutant property.
Given $h\in S$, it was proved by Deddens and Wong ([14]) that $\{M_{h}\}^{\prime}=\{M_{z}\}^{\prime}$. They obtained such result as a byproduct of a characterization of the commutant of an analytic Toeplitz operator $M_{\varphi}$ in terms of the commutants of the symbols of their inner-outer factorization. Such result was obtained later by Bourdon and Shapiro ([6]) as a byproduct of a deep result on intertwining analytic Toeplitz operators, which is intimately connected on a still open conjecture posed by Deddens.
In fact, such result can be obtained using the ideas of Shields and Wallen used in Theorem 2.1. For the sake of complementness we include a proof of this fact that uses only function theory.
Theorem 3.1 (Deddens-Wong).
Assume that $\varphi$ single cover a non-empty neighbourhood $W$ of $\varphi(\mathbb{D})$, then $\{M_{\varphi}\}^{\prime}=\{M_{z}\}^{\prime}$.
Proof.
To see the ideas clearly, let’s divide the proof into several steps.
Steep 1. Let $\varphi\in H^{\infty}(\mathbb{D})$, continuous in $\overline{\mathbb{D}}$. If for some $a\in\mathbb{D}$, $\varphi(z)-\varphi(a)$ has only a simple zero in $\mathbb{D}$ and $X$ commutes with $M_{\varphi}$ then $(Xf)(a)=(X1)(a)f(a)$.
Note that $\varphi(z)-\varphi(a)=(z-a)g(z)$, with $g$ continuous and non-vanishing in $\overline{\mathbb{D}}$. By the maximum principle $1/g\in H^{\infty}(\mathbb{D})$.
On the other hand, if $h\in H^{2}(\mathbb{D})$, and $h(a)=0$, then $h(z)=(z-a)u(z)$ for some $u\in H^{2}(\mathbb{D})$. Therefore,
$h(z)=\varphi(z)-\varphi(a)=u(z)\frac{u(z)}{g(z)}.$
Next, let $f\in H^{2}(\mathbb{D})$, then $f(z)-f(a)$ vanishes at $z=a$. Hence, it can be written as $f(z)-f(a)=(\varphi(z)-\varphi(a))F(z)$ for some $F\in H^{2}(\mathbb{D})$.
Since $X$ commutes with $M_{\varphi}$, we get
$X(f-f(a))(z)=X(M_{\varphi}F-\varphi(a)F)(z)=\varphi(z)XF(z)-\varphi(a)XF(z).$
Therefore, $0=X(f-f(a))(a)=Xf(a)-f(a)X1(a)$, which proves the first steep.
Steep 2. Let $\varphi\in H^{\infty}(\mathbb{D})$, such that $\varphi$ single covers a non-empty open subset $W\subset\varphi(\mathbb{D})$. If $X$ commutes with $M_{\varphi}$ then $Xf(a)=(X1)(a)f(a)$. The proof is the same as in Steep 1 as long we can write $\varphi(z)-\varphi(a)=(z-a)g(z)$; $z\in\mathbb{D}$ with $g$ and $1/g\in H^{\infty}(\mathbb{D})$.
Let $V=\varphi^{-1}(W)$, thus, $\varphi$ is univalent in $V$. Therefore, if $a\in V$, $\varphi^{\prime}(a)\neq 0$ and we can write:
$\varphi(z)-\varphi(a)=(z-a)g(z),\,\,z\in\mathbb{D}.$
Since $\varphi\in H^{\infty}(\mathbb{D})$, $g$ is bounded in $\mathbb{D}$. Moreover, by our hypothesis $\varphi(z)-\varphi(a)$ only vanish in $\mathbb{D}$ at the point $a$. Thus $g(z)\neq 0$ for all $z\in\mathbb{D}$. We claim that for all $z\in\mathbb{D}$, $|g(z)|>c$ for some $c>0$.
Let us consider a ball $B(a,r)$ centered at $a$ of radius $r>0$ with $\overline{B(a,r)}\subset V$. Then $\varphi^{-1}(\varphi(B(a,r))=B(a,r)$. Assume that $g$ is not bounded below in $\mathbb{D}$. Then, there is a sequence $(z_{n})\subset\mathbb{D}$ with $g(z_{n})\to 0$. Since $g(z)\neq 0$ for all $z\in\mathbb{D}$, $(z_{n})$ has to approach $\partial\mathbb{D}$. Thus, by choosing $n$ large enough, $z_{n}\notin B(a,r)$ and $\varphi(z_{n})\in\phi(B(a,r))$ which is a contradiction.
Steep 3. Finally, let us show that $\{M_{\varphi}\}^{\prime}=\{M_{z}\}^{\prime}$. Since $M_{z}$ has a minimal commutant then $\{M_{z}\}^{\prime}\subset\{M_{\varphi}\}^{\prime}$. Set $X\in\{M_{\varphi}\}^{\prime}$. Since the function $a\in\mathbb{D}\to k_{a}(z)$ is analytic we get that for each $f\in H^{2}(\mathbb{D})$ the function
$\langle Xf,k_{a}\rangle=h_{f}(a)$
is analytic on $\mathbb{D}$. Moreover, since $X$ is bounded by the Cauchy-Schwarz inequality we get that $h_{f}$ is bounded on $\mathbb{D}$. By steep 2 we get that for any $a\in W$ $h_{f}(a)=\langle Xf,a\rangle=(Xf)(a)=(X1)(a)f(a),$ In particular for $f=1$, we obtain that $h_{1}(z)=(X1)(z)\in H^{\infty}(\mathbb{D})$. Hence for all $a\in\mathbb{D}$, $Xf(a)=(X1)(a)f(a)$, that is $X$ is an analytic Toeplitz operator with symbol $(X1)(z)$ and therefore $X\in\{M_{z}\}^{\prime}$ as we desired. ∎
Let us shown that the elements in the class $S$ do not have the double commutant property.
Proposition 3.2.
Assume that $\varphi\in S$, that is, $\varphi$ is not univalent and $\varphi$ single cover a non-empty neighbourhood $W$ of $\varphi(\mathbb{D})$. Then $M_{\varphi}$ has not the double commutant property.
Proof.
Since $\varphi$ is not univalent then the polynomials on $\varphi$ are not dense in $H^{2}(\mathbb{D})$. Thus there exists $f_{0}\in H^{2}(\mathbb{D})\setminus\{0\}$ such that
$\langle p(M_{\varphi})1,f_{0}\rangle=0$
for each polynomial $p$. Let us consider the linear functional $\gimel:\mathcal{L}(H^{2}(\mathbb{D}))\to\mathbb{C}$ defined by:
$\gimel(X)=\langle X1,f_{0}\rangle.$
Since $H^{\infty}(\mathbb{D})$ is dense in $H^{2}(\mathbb{D})$ and $f_{0}\neq 0$, there exists $h\in H^{\infty}(\mathbb{D})$ such that $\langle h,f_{0}\rangle\neq 0$. Since $\varphi\in S$, by Deddens-Wong’s Theorem 3.1, we get that $M_{h}$ commutes with each element in $\{M_{\varphi}\}^{\prime}=\{M_{z}\}^{\prime}$, that is $M_{h}\in\{M_{\varphi}\}^{\prime\prime}$.
By continuity we get that $\gimel$ is a linear functional such that $\gimel(Y)=0$ for all $Y\in\overline{\textrm{Alg}(M_{\varphi}))}^{WOT}$ and $\gimel(M_{h})=\langle h,f_{0}\rangle\neq 0$. Thus, $M_{\varphi}$ don’t have the double commutant property as we desired to prove. ∎
Let us point out that T. R. Turner in his 1971 Michigan dissertation (see [29, Appendix D p.91]) posed the following question: Assume that $T$ has the double commutant property. Does $T^{n}$ have the double commutant property for all $n$? This question was solved by Deddens and Wogen in [13, Examples 1,2 p. 362]. Proposition 3.2 provides some more natural counterexamples for this question. Each function $\varphi\in S$ satisfies that $M_{\varphi}$ don’t have the double commutant property. The next example illustrate this fact.
Example 3.3.
If we consider the cardioid map $\varphi(z)=(z+1/2)^{2}$, in Figure 2 the region inside of the inner loop are double covered by $\varphi$. On the other hand, the region of the cardioid outside of the small loop is single covered by $\varphi$. Therefore $M_{\varphi}$ is an example of analytic Toeplitz operator that don’t have the double commutant property.
Moreover, if we consider the translation $\psi(z)=z+1/2$, by Walsh’s Theorem we get that $M_{\psi}$ has the minimal commutant property, therefore $M_{\psi}$ has trivially the double commutant property, but $M_{\psi}^{2}=M_{\varphi}$ does not.
Figure 2: The map $\varphi(z)=(z+1/2)^{2}$
4 The double commutant of $M_{z^{n}}$.
The commutant of $M_{z^{n}}$, $n\geq 2$ was studied by Z. Čučković in [31]. The commutant of $M_{z^{n}}$, $n\geq 2$ is very big. Čučković was able to describe the elements in the algebra of all Toeplitz operators that commutes with $M_{z^{n}}$. The commutant of $M_{z^{n}}$, $n\geq 2$ is so big, that the double commutant of $M_{z^{n}}$, $n\geq 2$ should be minimal.
In fact, this result is a special case of a general result by Turner [28] which asserts that if $X$ is a non-unitary isometry then $X$ has the double commutant property.
Nontheless, We wish to show an alternative proof of this fact for the special case $M_{z^{n}}$, $n\geq 2$.
Proposition 4.1.
The operator $M_{z^{n}}$, $n\geq 2$ has the double commutant property.
Proof.
Assume that $S$ is a linear operator in the double commutant of $M_{z^{n}}$. Since $S$ commutes with $M_{z}$ then $S=M_{h}$ for some bounded analytic function $h$.
For any analytic function $f(z)=\sum_{n=0}^{\infty}a_{n}z^{n}\in H^{2}(\mathbb{D})$, we can decompose
$f(z)=f_{1}(z)+zf_{2}(z)+\cdots+z^{n-1}f_{n}(z)$
where $f_{i}(z)=F_{i}(z^{n})$ for some $F_{i}\in H^{2}(\mathbb{D})$.
We consider the operator $Tf(z)=f_{1}(z)$ which is clearly bounded on $H^{2}(\mathbb{D})$ and $T$ commutes with $M_{z^{n}}$. Therefore forcing $M_{h}$ to commute with $T$ we get $M_{h}T1=h=h_{1}$, therefore $h=H_{1}(z^{n})$ for some $H_{1}\in H^{\infty}(\mathbb{D})$.
Next, the argument follows similar to the proof of Theorem 2.1 by Shields-Wallen. We consider $p_{n}$ the Féjèr polynomials of $H_{1}$ and we have that $p_{n}(z)\to H_{1}(z)$ uniformly on compact subsets of $\mathbb{D}$ and $\|p_{n}\|_{\infty}\leq\|H_{1}\|_{\infty}$. And this fact implies that $(p_{k}(M_{z^{n}}))$ converges to $M_{h}=M_{H_{1}(z^{n})}$ in the weak operator topology. ∎
To avoid do no repeat arguments, we state the following lemma whose proof follows by the end of the proof of Proposition 4.1.
Lemma 4.2.
Assume that $h\in H^{\infty}(\mathbb{D})$ and $B$ is an inner function. If there exists $\varphi\in H^{\infty}(\mathbb{D})$ such that $h(z)=\varphi(B)$ then $M_{h}\in\overline{\textrm{Alg}(M_{B})}^{WOT}$.
Remark 4.3.
The commutant of $M_{z^{n}}$ is so big, that there are elements which are not in the Toepliz algebra. For intance, set $a=re^{i\theta_{0}}\in\mathbb{D}$ and let us consider $a_{k}=r^{1/n}e^{\frac{\theta_{0}+2k\pi}{n}i}$, $k=0,\cdots,n-1$ the $n$-roots of $a$. Set $\lambda_{0}=e^{2\pi i/n}\in\partial\mathbb{D}$ the “first” root of the unity. Clearly $a_{k}=a_{0}\lambda_{0}^{k}$, for $k=0,\cdots,n-1$. The dilation operator defined by $L_{\lambda_{0}}f(z)=f(\overline{\lambda_{0}}z)$, also commutes with $M_{z^{n}}$. If we force $M_{h}$ to commute with $L_{\lambda_{0}}^{k}$, $k=1,...,n$ we get that $h$ is constant on each $z^{n}$-fiber, $(z^{n})^{-1}$, therefore there is a unique analytic funtion $\varphi$ such that $h(z)=\varphi(z^{n})$, and from this, we get again that $M_{z^{n}}$ has the double commutant property.
For future references, we can deduce the following corollary:
Corollary 4.4.
If $X\in\{M_{z^{k}}\}^{\prime\prime}$ then there exist $h\in H^{\infty}(\mathbb{D})$ such that $X=M_{h(z^{k})}$.
For the sake of complement the exposition we will extend the idea of Proposition 4.1 for any inner function $g$. For any isometry $V$ on a Hilbert space $H$, Wold decomposition theorem asserts that $H=H_{0}\oplus\sum_{n=1}^{\infty}\oplus V^{n}(H_{1})$, where $H_{1}=H\ominus V(H)$ and $H_{0}=\bigcap_{n=1}^{\infty}V^{n}(H)$. When $H=H^{2}(\mathbb{D})$ and $V$ is multiplication $M_{g}$ by a inner function $g$, then $H_{0}=\{0\}$ and
$H^{2}(\mathbb{D})=\sum_{n=1}^{\infty}\oplus s_{n}(z)H^{2}[g]$
when $H^{2}[g]=\{f(g)\,:\,f\in H^{2}(\mathbb{D})\}$ and $\{s_{1}(z),\cdots,s_{n}(z)\,\cdots\}$ is an orthonormal basis of $H^{2}(\mathbb{D})\ominus gH^{2}(\mathbb{D})$ that can be constructed by the Gram–Schmidt process. If $g(0)=0$ we can take $s_{1}=1$ and if not, we can take $s_{1}(z)=K_{a}(z)$, where $K_{a}$ denotes the normalized reproducing kernel at $a$, and $g(a)=0$.
Proposition 4.5.
If $g$ is an inner function then $M_{g}$ has the double commutant property.
Proof.
If $X\in\{M_{g}\}^{\prime\prime}$ then $X=M_{h}$, $h\in H^{\infty}(\mathbb{D})$. It is sufficient to show that $h(z)=H(g)$ for some $H\in H^{\infty}(\mathbb{D})$. If $f\in H^{2}(\mathbb{D})$ then $f(z)=\sum_{n=1}^{\infty}s_{n}(z)F_{n}(g)$ for $F_{n}\in H^{2}(\mathbb{D})$ $n\geq 1$. And $s_{1}(z)$ is $1$ or $K_{a}(z)$. We consider the projection operator $Tf=s_{1}(z)F_{1}(g)$ which is bounded on $H^{2}(\mathbb{D})$ (see [20]) and commutes with $M_{g}$. Therefore forcing $M_{h}$ to commute with $T$ we get
$M_{h}Ts_{1}(z)=s_{1}(z)h(z)=T(hs_{1}(z))=s_{1}(z)H_{1}(g).$
Therefore $h=H_{1}(g)$. Since $g$ is inner and $h\in H^{\infty}(\mathbb{D})$ we get that $H_{1}\in H^{\infty}(\mathbb{D})$ as we desired to prove. ∎
For future references, we can deduce the following corollary:
Corollary 4.6.
Assume that $B$ is an inner function. If $X\in\{M_{B}\}^{\prime\prime}$ then there exist $h\in H^{\infty}(\mathbb{D})$ such that $X=M_{h(B)}$.
5 Double commutant property for analytic Toeplitz operators induced by entire functions and functions in the Thomson-Cowen class.
In this section we will explore when an analytic Toeplitz operator $M_{\varphi}$ induced by an entire function $\varphi$ or a function in the Thomson-Cowen’s class have the double commutant property. Although we will do a unified study of both classes, to motivate the reader, we are going to look at some examples.
Our double commutant property for analytic Toeplitz operators induced by entire functions uses a result by Baker, Deddens and Ullman [3].
[…]
$k(\varphi)=\inf\{n(\varphi(\gamma),a)\,:\,n(\varphi(\gamma),a)\neq 0\}.$
Theorem 5.1 (Baker-Deddens-Ullman).
If $\varphi$ is a non-constant entire function and $k=k(\varphi)$ then there exists an entire function $h$ such that $\varphi(z)=h(z^{k})$ and $k(h)=1$.
As byproduct of this result, they obtained that if $\varphi(z)=h(z^{k})$ then $\{M_{\varphi}\}^{\prime}=\{M_{z^{k}}\}^{\prime}$. Let us see the following geometric necessary condition, if the points of the region $\varphi(\mathbb{D})$ have different winding numbers with respect to the curve $\varphi(\gamma)$ then $M_{\varphi}$ don’t have the double commutant property.
Proposition 5.2.
Assume that $\varphi$ is an entire function with $k(\varphi)=k$. If there exists of a point $c\in\varphi(\mathbb{D})$ such that $n(\varphi(\gamma),c)=p$ and $p>k=k(\varphi)$ then $M_{\varphi}$ does not have the double commutant property.
Proof.
By Baker-Deddens-Ullman’s result, there is an open subset $W\subset\varphi(\mathbb{D})$ such that $n(\varphi(\gamma),a)=k=k(\varphi)$ for all $a\in W$, and $\varphi(z)=h(z^{k})$ for some entire function $h$ which satisfies that $n(h(\gamma),a)=1$ for any $a\in W$.
By the Argument Principle (see [1, p. 152]), we obtain that $h$ single covers the open subset $W$. Moreover, since $n(\varphi(\gamma),c)=p>k$ we obtain that $h$ is not univalent. Hence, by results of Section 3, we proved that $M_{h}$ don’t have the double commutant property. Specifically, since the polynomials on $h$ are not dense in $H^{2}(\mathbb{D})$ there exists $f\in H^{2}(\mathbb{D})$ such that for any polynomial $p$, $\langle p(h),f\rangle=0$, and since $H^{\infty}(\mathbb{D})$ is dense in $H^{2}(\mathbb{D})$ there exists $g\in H^{\infty}(\mathbb{D})$ such that $\langle g,f\rangle\neq 0$.
As a consequence we get that $\langle p(h(z^{k})),f(z^{k})\rangle=0$ and $\langle g(z^{k}),f(z^{k})\rangle\neq 0$. Denoting $\gimel(X)=\langle X1,f\rangle$, this linear functional in $\mathcal{L}(H^{2}(\mathbb{D}))$ satisfies:
$\gimel(p(M_{\varphi}))=\langle p(M_{h(z^{k})})1,f(z^{k})\rangle=0$ (3)
and
$\gimel(M_{g(z^{k})})=\langle M_{g(z^{k})}1,f(z^{k})\rangle\neq 0$ (4)
for any polynomial $p$. We claim that $M_{g(z^{k})}\in\{M_{\varphi}\}^{\prime\prime}$. In such a case, equations (3) and (4) proves that the operator $M_{\varphi}$ don’t have the double commutant property.
Indeed, by Baker-Deddens-Ullman’s result we know that $\{M_{\varphi}\}^{\prime}=\{M_{z^{k}}\}^{\prime}$, therefore $\{M_{\varphi}\}^{\prime\prime}=\{M_{z^{k}}\}^{\prime\prime}$. By Corollary 4.2 we obtain that $M_{g(z^{k})}\in\overline{\textrm{Alg}(M_{z^{k}})}^{WOT}$, and since $M_{z^{k}}$ has the double commutant property we get that $M_{g(z^{k})}\in\{M_{z^{k}}\}^{\prime\prime}=\{M_{\varphi}\}^{\prime\prime}$ as we desired. Therefore we have shown that $M_{\varphi}$ don’t have the double commutant property. ∎
By the Argument Principle, we can deduce from Proposition 5.2 the following special case included in Corollary 5.3. Nonetheless, instructively, we include an alternative proof that uses the pingeonhole principle.
Corollary 5.3.
Assume that $\varphi$ is an entire function with $k(\varphi)=k$. If the image of a point under $\varphi$ has $p>k=k(\varphi)$ preimages then $M_{\varphi}$ does not have the double commutant property.
Proof.
Assume that the complex number $c\in\varphi(\mathbb{D})\subset\mathbb{C}$ has $p$ pre-images ($p>k$), that is, there exist $a_{1},\cdots,a_{p}\in\mathbb{D}$ such that $\varphi(a_{1})=\varphi(a_{2})=\cdots=\varphi(a_{p})=c$. Since $M_{\varphi}-cI=M_{\varphi-c}$ and the commutant is invariant under translations by the identity, we can suppose without loss that $\varphi(a_{1})=\cdots=\varphi(a_{p})=0$.
Since $\varphi$ is an entire function, by Baker-Deddens- Ullman’s result we can deduce that $\{M_{\varphi}\}^{\prime}=\{M_{z^{k}}\}^{\prime}$ for $k=k(\varphi)$. Therefore, $\{M_{\varphi}\}^{\prime\prime}=\{M_{z^{k}}\}^{\prime\prime}$. In particular, $M_{z^{k}}\in\{M_{\varphi}\}^{\prime\prime}$.
We claim that $M_{z^{k}}\notin\overline{\textrm{Alg}(M_{\varphi})}^{WOT}$. Assume by contradiction that there exists a net $p_{d}(\cdot)$, $d\in D$ such that for any $h,g\in H^{2}$:
$\lim_{d}\langle p_{d}(M_{\varphi})^{\star}h,g\rangle=\langle M_{z^{n}}^{\star}h,g\rangle.$
Set $k_{a}(z)=1/(1-\overline{a}z)$ the reproducing kernel at $a\in\mathbb{D}$. Since $k_{a_{j}}\in\textrm{Ker}(M_{\varphi}^{\star})$ for $j=1,\cdots,p$, we get:
$\lim_{d}\langle p_{d}(M_{\varphi})^{\star}(k_{a_{j}}),1\rangle=\lim_{d}p_{d}(0)=\langle M_{z^{k(\varphi)}}^{\star}(k_{a_{j}}),1\rangle=\overline{a_{j}}^{k(\varphi)}.$
But, since $p>k(\varphi)\geq 1$, by using the $k(\varphi)$-valence of $z^{k(\varphi)}$ and the pingeonhole principle, there exist at least two values $1\leq i_{0}<j_{0}\leq p$ such that $a_{i_{0}}^{k(\varphi)}\neq a_{j_{0}}^{k(\varphi)}$, a contradiction. ∎
The equality $\varphi(z)=h(z^{k})$ in the Baker-Deddens- Ullman result suggests that the double commutant property for an analytic Toeplitz operator induced for an entire function, will be closely related to the approximation of the polynomials on $h(z^{k})$. But, in this case this approximation cannot be in $H^{\infty}(\mathbb{D})$ because such functions are not univalent. This fact suggest to consider small spaces. Indeed, given $g$ an inner function let us denote
$H^{\infty}[g]=\{f\circ g\,:\,f\in H^{\infty}(\mathbb{D})\}.$
We say that the polynomials of $\varphi(g)$ are weak-star dense in $H^{\infty}[g]$ if the weak-star closure of polynomials on $\varphi(g)$ contains the space $H^{\infty}[g]$, and in such a case we will say that $\varphi(g)$ is a weak-star generator on $H^{\infty}[g]$.
Lemma 5.4.
Assume that $g$ is an inner function and let us suppose that $g(\mathbb{D})=\mathbb{D}$. The function $\varphi$ is a weak star generator of $H^{\infty}(\mathbb{D})$ if and only if $\varphi(g)$ is a weak-star generator of $H^{\infty}[g]$.
Proof.
We denote
$\mathcal{P}=\{p(\varphi)\,:\,p\,\textrm{polynomial}\}$
and
$\mathcal{P}[g]=\{p(\varphi(g))\,:\,p\,\textrm{polynomial}\}.$
For $\alpha=1$, $h\in\mathcal{P}_{*}^{(1)}$ if and only if there exists a sequence of polynomials $p_{n}$ such that
$\|p_{n}(\varphi)\|_{\infty}\leq M\quad\textrm{and}\quad p_{n}(\varphi(a))\to h(a)\,\,\textrm{for each}\,a\in\mathbb{D},$ (5)
which implies that $\|p_{n}(\varphi(g))\|_{\infty}\leq M$ and $p_{n}(\varphi(g(a))\to h(g(a))$ for each $a\in\mathbb{D}$, therefore $h\in\mathcal{P}_{*}^{(1)}[g]$.
Conversely, if $\|p_{n}(\varphi(g))\|_{\infty}\leq M$ and $p_{n}(\varphi(g(a))\to h(g(a))$ for each $a\in\mathbb{D}$ then $\|p_{n}(\varphi)\|_{\infty}\leq M$ and $p_{n}(\varphi(b))\to h(b)$ for each $b\in g(\mathbb{D})=\mathbb{D}$. That is, $h\in\mathcal{P_{*}^{(1)}}$.
The result follows because inductively we get that for any ordinal $\alpha$ we have $h\in\mathcal{P}_{*}^{(\alpha)}$ if and only if $h(g)\in\mathcal{P}_{*}^{(\alpha)}[g]$. ∎
Let $\varphi\in H^{\infty}(\mathbb{D})$, we say that $\varphi$ is in the Thomson-Cowen’s class ($\mathcal{TC}(\mathbb{D})$) if there exists a point $\lambda\in\mathbb{D}$ such that the inner part of $\varphi-\varphi(\lambda)$ is a finite Blaschke product. It is well know that the class $\mathcal{TC}(\mathbb{D})$ contains all non-constant functions in $H^{\infty}(\overline{\mathbb{D}})$. Below we state the remarkable theorem due to Thomson and Cowen ([9, 27, 26]):
Theorem 5.5 (Thomson(1976)-Cowen(1978)).
Assume $\phi\in\mathcal{TC}(\mathbb{D})$. Then there exists a finite Blaschke product $B$ and a function $\psi\in H^{\infty}(\mathbb{D})$ such that $\phi=\psi(B)$ and $\{M_{\phi}\}^{\prime}=\{M_{B}\}^{\prime}$.
The following result characterize when an element in the Thomson-Cowen class, has the double commutant property.
Theorem 5.6.
Assume that $\varphi\in\mathcal{TC}(\mathbb{D})$ and set $\varphi=h(B)$ a Thomson-Cowen factorization of $\varphi$. The following conditions are equivalents
  1. 1.
    $M_{\varphi}$ has the double commutant property.
  2. 2.
    The polynomials on $\varphi$ are weak-star dense in $H^{\infty}[B]$.
Proof.
By way of contradiction let us suppose that $M_{\varphi}$ has the double commutant property and that the polynomials on $h(B)$ are not weak-star dense in $H^{\infty}[B]$. As in the proof of Theorem 2.6 (see Remark 2.7) we get that the polynomials on $h(B)$ are not WOT dense in $H^{\infty}[B]$. Hence, there exist functions $f\in H^{\infty}(\mathbb{D})$ and $\psi(B)\in H^{\infty}[B]$ such that for any polynomial $p$
[…]
and
[…]
Let us consider the linear functional $\gimel$ defined on the subspace $\{M_{g}\,:\,g\in H^{\infty}(\mathbb{D})\}$ as
[…]
and the sublinear function $p(T)=\|T\|$. Let us denote $\widetilde{\gimel}$ the Hahn-Banach’s sublinear extension of $\gimel$. By (6) and (7) we have that ~$\widetilde{\gimel}(M_{p\circ h(B)})=\widetilde{\gimel}(M_{p(\varphi)})=0$ for any polynomial $p$ and ~$\widetilde{\gimel}(M_{\psi(B)})\neq 0$. On the other hand, by Corollary 4.2 we get that $M_{\psi(B)}\in\{M_{B}\}^{\prime\prime}=\{M_{\varphi}\}^{\prime\prime}$, that is, we arrive to a contradiction, because by hypothesis $M_{\varphi}$ has the double commutant property.
Assume that the polynomials on $\varphi$ are weak-star dense in $H^{\infty}[B]$. Set $X\in\{M_{\varphi}\}^{\prime\prime}$. Since $\varphi=h(B)$, we have that $\{M_{B}\}^{\prime\prime}=\{M_{\varphi}\}^{\prime\prime}$.Therefore, by Corollary 4.6 we get that $X=M_{\psi(B)}$ for some $\psi\in H^{\infty}(\mathbb{D})$.
Indeed, let us show that for any ordinal $\alpha$
$\{M_{g}\,:\,g\in\mathcal{P}_{*}^{(\alpha)}[B]\}\subset\overline{\{M_{p(\varphi)}\,p\,\textrm{polynomial}\}}^{WOT}:=\mathcal{M}.$
Indeed for $\alpha=1$ if $g\in\mathcal{P}_{*}^{(1)}[B]$ there exists a sequence of polynomials $p_{n}$ such that $\|p_{n}(h(B))\|_{\infty}\leq C$ and $p_{n}(h(B(a)))\to g(a)$ for any $a\in\mathbb{D}$. By Remark 2.5 we get that $M_{p_{n}(h(B))}$ converges in the strong operator topology to $M_{g}$, therefore $M_{g}\in\mathcal{M}$. Now, let us assume that for an ordinal $\alpha$ we have $\{M_{g}\,:\,g\in\mathcal{P}^{(\alpha)}[B]\}\subset\mathcal{M}$. If $g\in\mathcal{P}_{*}^{(\alpha+1)}$ there exists a sequence of function $g_{l}\in\mathcal{P}_{*}^{(\alpha)}[B]$, such that $\|g_{l}\|_{\infty}<C$ and $g_{l}(a)\to g(a)$ for any $a\in\mathbb{D}$. By Remark 2.4 we have that $M_{g_{l}}\to M_{g}$ in the strong operator topology and $M_{g_{l}}\in\mathcal{M}$ for all $l$, since $\mathcal{M}$ is closed in the weak operator topology we get that $M_{g}\in\mathcal{M}$.
Since there exists an ordinal $\beta$ such that $\mathcal{P}_{*}^{(\beta)}[B]\supset H^{\infty}[B]$. We get that
$\{M_{g(B)}\,:\,g\in H^{\infty}(\mathbb{D})\}\subset\mathcal{M}=\overline{\{M_{p(\varphi)}\,:\,p\,\textrm{polynomial}\}}^{WOT}.$
That is, $X=M_{\psi(B)}\in\overline{\{M_{p(\varphi)}\,:\,p\,\textrm{polynomial}\}}^{WOT}$ as we desired to prove. ∎
Remark 5.7.
If we consider $\varphi(z)=z^{6}$ and $f(z)=z^{2}$, clearly $\varphi(z)=f(z^{3})$. Of course $\varphi(z)=z^{6}$ has the double commutant property but the polynomials on $\varphi$ are not weak-star dense in $H^{\infty}[z^{3}]$ and the polynomials on $z^{2}$ are not weak-star dense in $H^{\infty}(\mathbb{D})$. At the first glance one may think that something is wrong. The problem here is that the respresentation $\varphi(z)=z^{6}=f(z^{3})$ is not the Thomson-Cowen’s representation of $\varphi$, because the commutant of $M_{\varphi}$ is bigger than the commutant of $M_{z^{3}}$. In this case the representation of $z^{6}$ is just $z^{6}=f(z^{6})$ when $f(z)=z$. That is, equality of the commutants $\{M_{\varphi}\}^{\prime}=\{M_{B}\}^{\prime}$ in the Thomson-Cowen’s Theorem is fundamental.
Corollary 5.8.
Assume that $\varphi\in\mathcal{TC}(\mathbb{D})$, and $\varphi(z)=h(B)$ with $B$ a finite Blaskhe product. The following conditions are equivalents:
  1. 1.
    $M_{\varphi}$ has the double commutant property.
  2. 2.
    $M_{h}$ has the minimal commutant property.
Now, let us obtain the geometric result for functions in the Thomson-Cowen class. It is well know that the factorization $\varphi=h(B)$, is unique in the sense of modulo Moebius maps. We denote by $b(\varphi)$ the Thomson-Cowen order of $\varphi$, that is, the maximal order of finite Blaskhe product $B^{\prime}$ for which $\varphi=h^{\prime}(B^{\prime})$ for some $h^{\prime}\in H^{\infty}(\mathbb{D})$. For maps $\varphi\in H^{\infty}(\overline{\mathbb{D}})$ it is defined the minimal winding number as
$k(\varphi)=\inf\{n(\varphi(\gamma),\varphi(a)\,:\,\varphi(a)\notin\varphi(\partial\mathbb{D})\}.$
The map $\varphi\in H^{\infty}(\overline{\mathbb{D}})$ is said to have the Minimal Winding number property if $k(\varphi)=b(\varphi)$. Clearly, by Baker-Deddens and Ullmann’s result if $\varphi$ is an entire function then $\varphi$ has the Minimal Winding number property.
The following result is the analogous result to Proposition 5.2 for functions in $H^{\infty}(\overline{\mathbb{D}})$.
Theorem 5.9.
Assume that $\varphi\in H^{\infty}(\overline{\mathbb{D}})$. If there are two points $a,b\in\mathbb{D}$ such that $\varphi(a),\varphi(b)\notin\varphi(\partial\mathbb{D})$ and $n(\varphi(\gamma),\varphi(a))\neq n(\varphi(\gamma),\varphi(b))$, then $M_{\varphi}$ has not the double commutant property.
Proof.
Indeed, we can suppose without loss that $\varphi$ is not constant, in such a case $\varphi\in\mathcal{TC}(\mathbb{D})$. Set $\varphi(z)=h(B)$ a Thomson-Cowen factorization. If $M_{\varphi}$ has the double commutant property, by Theorem 5.6 the polynomials on $h$ are weak star dense in $H^{\infty}(\mathbb{D})$ therefore by [22, Proposition 3] we get that $h$ is univalent. Therefore, for any $a\in\mathbb{D}$ such that $\varphi(a)\notin\varphi(\partial\mathbb{D})$ we have that $n(\varphi(\gamma),\varphi(a))=b(B)$, a contradiction. ∎
Corollary 5.10.
Assume that $\varphi\in H^{\infty}(\overline{\mathbb{D}})$. If $M_{\varphi}$ has the double commutant property then $\varphi$ has the Minimal Winding number property.
Corollary 5.11.
Assume that $\varphi\in H^{\infty}(\overline{\mathbb{D}})$ and that $\varphi$ acting on $\partial\mathbb{D}$ is a Jordan curve. Then, $M_{\varphi}$ has the double commutant property if and only if $n(\varphi(\gamma),\varphi(a))$ is constant for each $a\in\mathbb{D}$ such that $\varphi(a)\notin\varphi(\partial\mathbb{D})$.
Proof.
The necessity part follows from Theorem 5.9. Now, we assume that $\varphi(z)=f(B)$. By Theorem 5.6 we get that $f$ is univalent. Since the boundary of $f(\mathbb{D})$ is a Jordan curve, we get that the polynomials on $f$ are weak-star dense in $H^{\infty}(\mathbb{D})$, therefore by Walsh Theorem $M_{\varphi}$ has the minimal commutant property. By Theorem 5.6 again, we get that $M_{\varphi}$ has the double commutant property as we wanted. ∎
The following example shows that the geometric condition in Theorem 5.9, in general, is not sufficient to guarantee the double commutant property of an analytic Toeplitz operator.
Example 5.12.
Let $f$ be a univalent function which maps $\mathbb{D}$ onto the slit disk $\mathbb{D}\setminus[0,1)$. If we consider $\varphi(z)=f(z^{p})$ we have that for any $c\in\varphi(\mathbb{D})$, $M_{\varphi}$ don’t have the double commutant property.
6 Concluding remarks and open questions
In [10] the authors ask if there exists a function $h\in H^{\infty}(\mathbb{D})$ that is not in the Thomson-Cowen’s class. A negative answer to this question provides that the results included in this paper could fully characterize when an analytic Toeplitz operator has the double commutant property.
We dont know if the geometric condition in terms of the winding number is also sufficient for the double commutant property in some cases. More specifically, assume that $\varphi$ is an entire function $\varphi$ and let us suppose that for any $a\in\mathbb{D}$ $n(\varphi(\gamma),\phi(a))=1$. It is $\varphi$ univalent on $\partial\mathbb{D}$?
Funding information. M.J.G. was supported in part by the Spanish Ministerio de Ciencia e Innovación (grant no. PID2021-123151NB-I00). This publication is part of the project PID2022-139449NB-I00, funded by MCIN/AEI/10.13039/501100011033/FEDER, UE. The authors were supported by Grant “Operator Theory: an interdisciplinary approach,” reference ProyExcel__\__00780, a project financed in the 2021 call for Grants for Excellence Projects, under a competitive biddingregime, aimed at entities qualified as Agents of the Andalusian Knowledge System, in the scope of the Andalusian Research, Development and Innovation Plan (PAIDI 2020). Counseling of University, Research and Innovation of the Junta de Andalucía.